OpenStudy (anonymous):

given exact and approximate solutoins to three decimal places y^2-20y+100=4 would the answer then be y=11???

OpenStudy (anonymous):

no 11 does not work first of all there should probably be two answers. do you want to factor?

OpenStudy (anonymous):

try \[(x-10)^2=4\] so \[x-10=2\] or \[x-10=-2\] then solve each equation separately

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

i think i was working it out as a reg, equation

OpenStudy (anonymous):

it would be 12 right

OpenStudy (anonymous):

one would be 8

OpenStudy (anonymous):

for an equation with an exponent of 2 you have to factor or use the quadratic formula. yes 12 is one answer but 8 is also an answer because if you solve \[x-10=-2\] you get \[x=8\]

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

cool i got it right thank you ,